QUESTION 10 Currently the Australian Government is aiming to reduce Australias carbon emissions to 5% below 2000 emissions levels by 2020 Reforms Emissions change (compared to 2000 | 10% increase levels) by 2020 Total cost forindustries to introduce 0 reform measures (billions of dollars No change 15% 10% reduction reduction 0.3 a. If the Government introduced a permit system, resulting in total industry costs being only $100 million, what would be the expected 2020 pollutions levels compared to 2000 levels? Equal to or GT for Greater Than Type LT for Less Than, E for b. Consider the agricultural sector is responsible for 16% of Australias carbon emissions. The Government decides to introduce the reform option C. How much would it cost the agricultural sector as a result of introducing this reform? Assume the cost is directly proportionally to all industries within the Australian economy. Answer to two decimal places in billions of dollars. $ billion C. Consider a biscuit company decides to contribute toward the reduction in carbon emissions. It does so by purchasing $10 million worth of permits issued by the government. The company then spends another $8 million on upgrading its manufacturing processes. What is the net cost to society resulting from these decisions? Answer to the nearest whole number in millions of dollars (with no decimal places). $ million
Answer #1

Ans ) a) If govt introduced a permit system then the expected pollution level by 2020 will be reduction by 5% in the total emission. Or LT (less than) 2000 emission.

Ans )b) the agricultural sector pollution= 16%

According to question option c gives 5% reduction with the cost of 1.2 billion dollars then:

1.2=5%

1.2*100/5= 1%

1%= 240

And 16%= 16*240

= 3840( billion dollars)

Ans )c) 54(billion dollars)

1.2+ 1.5+ 0.3 = 3.0 %

3/100*18

= 54( billion dollars)